OpenStudy (anonymous):

Hyperbola Explanation Find the center, vertices, foci, eccentricity, and asymptotes of the hyperbola with the given equation:

OpenStudy (anonymous):

\[\frac{ y^{2} }{ 25 }-\frac{ x^{2} }{ 144 } = 1\] This is a practice problem from a website, and I'd like to know how to do all of this.

OpenStudy (anonymous):

@hero @ganeshie8 @dan815

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@hartnn

ganeshie8 (ganeshie8):

see if this helps

OpenStudy (dan815):

okay hmmm

OpenStudy (dan815):

do a hyperbola have the foci in the same place as the ellipse

OpenStudy (anonymous):

Okay, so: Vertical No x intercepts Y intercept: (0,-+25) Vertices: (0,-+25) LTA: 50 LCA: 288 Foci: 13 e > 1 Asymptotes: y = -+25/144x

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

idk dan

OpenStudy (dan815):

ah ganeshie8's post cover it all

OpenStudy (anonymous):

Pretty cool chart, makes it easy. :D

OpenStudy (anonymous):

Give ganeshie a medal please.

OpenStudy (dan815):

=]

OpenStudy (dan815):

i give ganehie8 too many medals :)

OpenStudy (anonymous):

lol

ganeshie8 (ganeshie8):

\[\frac{ y^{2} }{ 25 }-\frac{ x^{2} }{ 144 } = 1\]

ganeshie8 (ganeshie8):

\[\frac{ y^{2} }{ 5^2 }-\frac{ x^{2} }{ 12^2 } = 1\]

ganeshie8 (ganeshie8):

\(\large a = 5\), \(\large b = 12\)

ganeshie8 (ganeshie8):

\(\large c = \sqrt{a^2+b^2} = \sqrt{5^2 + 12^2} = 13\)

ganeshie8 (ganeshie8):

use the chart now :)

OpenStudy (anonymous):

Oh yeah, I have to square root them to get the actual equation to use.